OpenStudy (anonymous):

A girl coasts down a hill on a sled, reaching a level surface at the bottom with a speed of 7.6 m/s. The coefficient of kinetic friction between runners and snow is 0.052, and the girl and sled together weigh 744 N. The acceleration of gravity is 9.8 m/s 2 . How far does the sled travel on the level surface before coming to a rest? Answer in units of m

OpenStudy (anonymous):

Ff = .052*744N a= Ff/m = Ff/(744N/g) a = .052g use in: Vf^2 = 0 = Vi^2 -2*a*d to find d